# Misreading EPR: Variations on an Incorrect Theme

### v3

@@ 1 Introduction
Some of the examples we will see come from popularizations of science. Accordingly, in Section 2 we will go through the EPR thought-experiment in the manner, one might say, of a publication that has “New York” in its name. All of the errors we will put on display can be appreciated with that background. Bits of mathematical notation in the more “serious” exhibits may make their prose harder to follow, but the formulae are only a smokescreen over the conceptual problems.
@@ 2 EPR, Very Quickly
The EPR thought-experiment rests upon the fact that in quantum physics, it is possible to prepare a pair of objects in such a way that, once we perform a measurement of our choice on one object, we can then make a completely confident prediction about the outcome of a corresponding measurement on the other. We make a choice and take action here, and then we can foretell what will happen when we act over there. The original EPR presentation considered position and momentum as the possible “observables”: Of each particle, we can ask where it is, or alternatively, how it is moving.
@@
It was important for EPR that the second particle be far away, because then they thought they could safely assume that it cannot be influenced by what we choose to do on the first particle. Schrödinger later used the analogy of a student getting quizzed in an oral examination [4]. The student is too tired to answer more than one question, so they will get the first question correct, but their answer to the second will be pretty much random. However, because they do not know ahead of time what the first question will be, they have to know all the answers going in. The distant particle is like this fatigued student. It has to “know” both the position and the momentum of the nearby particle, even though in any specific experiment it can reveal only one of the two. This amounts to saying that it must have both a position and a momentum before we measure it.
@@
Other people came to this topic after Bohm, and we will discuss some of their contributions later. First, we need to clear the air of misconceptions about what EPR said.
@@ 3 Becker
There is no choice between alternatives in Becker’s description, just a single procedure to follow. EPR recognized that making a selection was a significant step, but Bohr took that point further. Much of Bohr’s reply to EPR was devoted to devising a more specific implementation scheme for their thought-experiment [8]. This mattered to Bohr, because he wished to emphasize that the choice of measuring either position or momentum requires selecting one of two mutually exclusive laboratory setups [9, pp. 195–6]. We will return to the topic of Bohr’s reply later (§12), but for now, we should underline the fact that missing the ingredient of choosing between experimental arrangements means that this portrayal fails to engage with a crucial part of Bohr’s thinking. This is particularly important because Becker’s book is in large part a polemic against Bohr’s viewpoint, or at least a viewpoint that it attributes to Bohr. Later printings fixed this error, but left others (we will touch on a systemic problem in §12). I bring this erratum to the forefront not just as a conversation piece about how quality control does or does not happen in pop science — although that is a conversation I believe we need to have! — but also because it provides a clean introduction to the other misreadings of EPR that we will survey as we go along.
@@ 6 Weinberg
Again, we have the measurements being applied in succession on the same object. This explicitly contradicts Bohm, who insists, “the investigator can measure either the $x$, $y$, or $z$ component of the spin of particle No. 1, but not more than one of these components, in any one experiment” [7, p. 614]. The final sentence adds to the confusion, as it appears to say that the second measurement, the one on the $x$-component, will force the previously established value of the $z$-component into remission. In addition to missing the point that the first measurement consumes the entanglement, this line also confounds the EPR criterion of reality, which is based on the assumption that the measurement of one particle does not disturb the other.
@@ 10 New Scientist and Carroll
No, it isn’t. It’s an unremarkable possibility familiar from everyday life. The entire buildup to this declaration is beside the point. [^5] Part of that buildup includes an assertion that a $\psi$ function must be “real” and “not just a bookkeeping device to keep track of probabilities” (p. 237). This claim is unfounded; Carroll’s brief argument for it neglects the fact that probabilities for different, mutually exclusive scenarios do not have to fit together in the way that classical physics assumes they do [26, 27]. This point was appreciated by Feynman in 1948 [28], and Born and Heisenberg seem to have had at least a finger on it as early as 1927 [29]. None of the conceptual or mathematical apparatus of quantum theory is necessary for Carroll’s scenario, and a big sign of why is that the story considers only one observable, the location, of each character.
@@ 12 Interlude: The Shorter Bohr
But that is exactly what Bohr did. A possible reply in the Bohrian vein: “Could a ‘reasonable definition of reality’ permit so basic a fact as the simultaneity of two events to be dependent on the observer’s frame of reference? Many notions familiar from everyday life only become well-defined in relativity theory once we fix a Lorentz frame. Likewise, many statements in quantum theory only become well-defined once we have given a complete description of the experimental apparatus and its arrangement.”
@@ 13 EPR and Beyond the Intrinsic
One way to make this way of thinking more precise is to ask how exactly quantum theory resists being fit into the phase-space picture. In classical mechanics, if we know the exact coordinates of a system in phase space, we can predict anything else. Perhaps surprisingly, quantum theory allows for a relaxed version of this: instead of a full-fledged phase space, an experiment with the property that, if one has statistics for the potential outcomes of it, one can calculate the probabilities for the possible outcomes of any other measurement [45, 46, 47]. Another way to express this is that, if one establishes an informationally complete measurement as a standard reference, then any quantum state is equivalent to a probability distribution. In the Bohm version of EPR, our systems of interest are “spins”; if a particle is “spin-$j$”, then an informationally complete measurement for that spin degree of freedom must have at least $(2j+1)^{2}$ outcomes [48, 49, 50]. The degree to which such a reference measurement can be made to resemble the classical ideal of reading off the phase-space coordinates is a measure of how far quantum theory deviates from classical physics [51, 52, 53]. Getting a clean and precise expression of this deviation requires finding the optimal reference measurements. To the surprise of the people who have worked on this topic, the question of finding the optimal reference measurements turns out to have intriguing connections with far-flung areas of mathematics [54, 55, 56, 57, 58, 59].
@@ Acknowledgments
I thank Chris Fuchs for telling me about the example from Becker’s book, and for suggesting a few other places I should look. Also, I thank N. David Mermin for bringing the Pagels example to my attention, and for helpful discussions (about the Pagels and Weinberg passages in particular).

### v4

@@ 1 Introduction
Some of the examples we will see come from popularizations of science. Accordingly, in Section 2 we will go through the EPR thought-experiment in the manner of a glossy pop-science magazine of yore. All of the errors we will put on display can be appreciated with that background. Bits of mathematical notation in the more “serious” exhibits may make their prose harder to follow, but the formulae are only a smokescreen over the conceptual problems.
@@ 2 EPR, Very Quickly
The EPR thought-experiment rests upon the fact that in quantum physics, it is possible to prepare a pair of objects in such a way that, once we perform a measurement of our choice on one object, we can then make a completely confident prediction about the outcome of a corresponding measurement on the other. We make a choice and take action here, and then we can foretell what will happen when we act over there. The original EPR presentation considered position and momentum as the possible “observables”: Of each particle, we can ask where it is, or alternatively, how it is moving. These are a pair of “observables” to which the uncertainty principle applies. According to the rules of quantum mechanics, there is a tradeoff of predictability between the two. For any single quantum particle, if we can make a precise prediction of a position measurement, we cannot make a precise prediction of a momentum measurement, and vice versa. This raises the question of how to interpret that tradeoff. Is it just that we are big and electrons are small, so our attempts to measure them are always clumsy and disrupt what they’re doing… Or is the truth something stranger? Does a quantum particle have a momentum before we go and measure it?
@@
It was important for EPR that the second particle be far away, because then they thought they could safely assume that it cannot be influenced by what we choose to do on the first particle. As Einstein said, in the course of developing his thought-experiments on the topic: “Thus, without any experiment on [a particle], it is possible to predict, according to a free choice, either the momentum or the position of [it] with in principle arbitrary accuracy” [4]. Schrödinger later used the analogy of a student getting quizzed in an oral examination [5]. The student is too tired to answer more than one question, so they will get the first question correct, but their answer to the second will be pretty much random. However, because they do not know ahead of time what the first question will be, they have to know all the answers going in. The distant particle is like this fatigued student. It has to “know” both the position and the momentum of the nearby particle, even though in any specific experiment it can reveal only one of the two. This amounts to saying that it must have both a position and a momentum before we measure it.
@@
One can carry the argument a step further. Suppose that you have one half of a Bohm-style pair of atoms, and you haven’t yet decided whether you want to do the H experiment or the V experiment. If you pick the H experiment, you will be able to make an H prediction, and so using the EPR criterion, you can say that the other atom has an H property. Likewise, if you pick the V experiment, you will be able to use the EPR criterion to say that the other atom has a V property. Before you do a measurement and have the result in hand, you can’t employ the EPR criterion to make a specific statement. For example, until you do a V measurement on the first atom, you can’t say what the value of the second atom’s V property is — whether it’s “swerve north” or “swerve south”. But you can say, ahead of time, that the atom must have a V property, and also an H property. You don’t know which value they have, but they’ve got to be there. And because they aren’t part of how quantum mechanics describes the situation, quantum mechanics must be incomplete — or so EPR would say.
We won’t need to worry about this extra wrinkle too much. The problems with many retellings of EPR can be seen without it. The idea of deducing that properties exist ahead of time without saying exactly what values they take really comes into its own in later arguments that cut into the quantum mysteries more deeply [9]. Other people came to this topic after Bohm, and we will discuss some of their contributions momentarily. First, we need to clear the air of misconceptions about what EPR said.
@@ 3 Becker
There is no choice between alternatives in Becker’s description, just a single procedure to follow. EPR recognized that making a selection was a significant step, but Bohr took that point further. Much of Bohr’s reply to EPR was devoted to devising a more specific implementation scheme for their thought-experiment [10]. This mattered to Bohr, because he wished to emphasize that the choice of measuring either position or momentum requires selecting one of two mutually exclusive laboratory setups [11, pp. 195–6]. We will return to the topic of Bohr’s reply later (§12), but for now, we should underline the fact that missing the ingredient of choosing between experimental arrangements means that this portrayal fails to engage with a crucial part of Bohr’s thinking. This is particularly important because Becker’s book is in large part a polemic against Bohr’s viewpoint, or at least a viewpoint that it attributes to Bohr. Later printings fixed this error, but left others (we will touch on a systemic problem in §12). I bring this erratum to the forefront not just as a conversation piece about how quality control does or does not happen in pop science — although that is a conversation I believe we need to have! — but also because it provides a clean introduction to the other misreadings of EPR that we will survey as we go along.
@@ 6 Weinberg
Again, we have the measurements being applied in succession on the same object. This explicitly contradicts Bohm, who insists, “the investigator can measure either the $x$, $y$, or $z$ component of the spin of particle No. 1, but not more than one of these components, in any one experiment” [8, p. 614]. The final sentence adds to the confusion, as it appears to say that the second measurement, the one on the $x$-component, will force the previously established value of the $z$-component into remission. In addition to missing the point that the first measurement consumes the entanglement, this line also confounds the EPR criterion of reality, which is based on the assumption that the measurement of one particle does not disturb the other.
@@ 10 New Scientist and Carroll
No, it isn’t. It’s an unremarkable possibility that could occur in everyday life. The entire buildup to this declaration is beside the point. [^5] Part of that buildup includes an assertion that a $\psi$ function must be “real” and “not just a bookkeeping device to keep track of probabilities” (p. 237). This claim is unfounded; Carroll’s brief argument for it neglects the fact that probabilities for different, mutually exclusive scenarios do not have to fit together in the way that classical physics assumes they do [28, 29]. This point was appreciated by Feynman in 1948 [30], and Born and Heisenberg seem to have had at least a finger on it as early as 1927 [31]. None of the conceptual or mathematical apparatus of quantum theory is necessary for Carroll’s scenario, and a big sign of why is that the story considers only one observable, the location, of each character.
@@ 12 Interlude: The Shorter Bohr
But that is exactly what Bohr did. A possible reply in the Bohrian vein: “Could a ‘reasonable definition of reality’ permit so basic a fact as the simultaneity of two events to be dependent on the observer’s frame of reference? Many notions familiar from everyday life only become well-defined in relativity theory once we fix a Lorentz frame. Likewise, many statements in quantum theory only become well-defined once we have given a complete description of the experimental apparatus and its arrangement. If you set up your laboratory to do a position measurement, you can make a position prediction. If you instead arrange your equipment to do a momentum measurement, you can make a momentum prediction. And before you take your equipment off the shelf, when all your devices are still in the cabinet, your laboratory is in the null configuration. Trying to blend deductions from multiple different, mutually exclusive arrangements will give a meaningless result, just like mixing up reference frames in relativity will make you blunder into what looks like a paradox.”
@@ 13 EPR and Beyond the Intrinsic
One way to make this way of thinking more precise is to ask how exactly quantum theory resists being fit into the phase-space picture. In classical mechanics, if we know the exact coordinates of a system in phase space, we can predict anything else. Perhaps surprisingly, quantum theory allows for a relaxed version of this: instead of a full-fledged phase space, an experiment with the property that, if one has statistics for the potential outcomes of it, one can calculate the probabilities for the possible outcomes of any other measurement [47, 48, 49]. Another way to express this is that, if we are sufficiently clever, we can establish a standard reference measurement. Having probabilities for all things that could happen when we apply the reference measurement to a system is mathematically the same as writing a quantum state for that system. Instead of talking about “$\psi$ functions” like in the old days, all we need are those probabilities. In the Bohm version of EPR, our systems of interest are “spins”; if a particle is “spin-$j$”, then a reference measurement for that spin degree of freedom must have at least $(2j+1)^{2}$ outcomes [50, 51, 52]. The degree to which such a reference measurement can be made to resemble the classical ideal of reading off the phase-space coordinates is a measure of how far quantum theory deviates from classical physics [53, 54, 55]. Getting a clean and precise expression of this deviation requires finding the optimal reference measurements. To the surprise of the people who have worked on this topic, the question of finding the optimal reference measurements turns out to have intriguing connections with far-flung areas of mathematics [56, 57, 58, 59, 60, 61].
Another line of attack uses that idea of knowing, even before you pick which measurement to do, that you can invoke the EPR criterion. Recall how the argument went: If you do the H measurement on one half of an EPR pair, you’ll be able to predict the result of an H measurement upon the other half. Likewise, if you do the V measurement on the first particle, you’ll then be able to predict the result of a V measurement upon the second particle. So, before you set up your magnet in either orientation, you know that you will be able to make some prediction. You can tell ahead of time that you will be able to employ the EPR criterion and say that the second particle has some definite value of a property. You just don’t yet know which value it will turn out to be. On the assumption that the second particle really is undisturbed (and also that the EPR criterion is good), you can already say that the second particle must have a V value and an H value.
Several people noticed that if you push this argument harder, it breaks down. If we consider not just two alternatives, like H versus V, but instead a well-chosen menu of possible experiments with more options, then the set of properties deduced for the second particle will be logically inconsistent with itself. Chasing through the equations, one finds that in some measurement, the particle is guaranteed to swerve north, and also guaranteed to swerve south. This inconsistency forces us back to our assumptions: Something we have presumed must be wrong [9].
@@ Acknowledgments
I thank Chris Fuchs for telling me about the example from Becker’s book, and for suggesting a few other places I should look. Also, I thank N. David Mermin for bringing the Pagels example to my attention, and for helpful discussions (about the Pagels and Weinberg passages in particular). Arthur Fine’s comments about v3 of this essay led me to post a v4.